OpenStudy (anonymous):

What is the x-coordinate of the x-intercept of the line that passes through points (-5, 3) and (1, -3)?

OpenStudy (jhannybean):

in order to solve this you need to find the slope first. \[\large m= \frac{ y_{2}-y_{1} }{ x_{2}-x_{1} }= \frac{ -3-(3) }{ 1-(-5) }\]\[m= \frac{ -6 }{ 6 }=-1\] now that we've found the slope we plug it into point-slope form, \[\large y-y_{1}=m(x-x_{1})\] with our slope (m) = -1 and our points (-5,3) and (1,-3) \[\large y-3=-1(x-(-5))\]\[\large y-3=-1(x+5)\]\[\large y-3 =-x-5\]\[\large y-3+3=-x-5+3\]\[\large y=-x-2\] Now we graph this line to find out x-coordinate of our x-intercept